# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v1

@@ Abstract
Post-hoc explanation methods for black-box models often struggle with faithfulness and human interpretability due to the lack of explainability in current neural models. Meanwhile, B-cos networks have been introduced to improve model explainability through architectural and computational adaptations, but their application has so far been limited to computer vision models and their associated training pipelines. In this work, we introduce B-cos LMs, i.e., B-cos networks empowered for NLP tasks. Our approach directly transforms pre-trained language models into B-cos LMs by combining B-cos conversion and task fine-tuning, improving efficiency compared to previous B-cos methods. Our automatic and human evaluation results demonstrate that B-cos LMs produce more faithful and human interpretable explanations than post hoc methods, while maintaining task performance comparable to conventional fine-tuning. Our in-depth analysis explores how B-cos LMs differ from conventionally fine-tuned models in their learning processes and explanation patterns. Finally, we provide practical guidelines for effectively building B-cos LMs based on our findings. Our code is available at https://anonymous.4open.science/r/bcos_lm.
B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability
Yifan Wang1,3,  Sukrut Rao2,3,  Ji-Ung Lee1,3, Mayank Jobanputra1  and Vera Demberg1,2,3 1Saarland University, Saarbrücken, Germany 2Max Planck Institute for Informatics, Saarland Informatics Campus, Saarbrücken, Germany 3RTG Neuroexplicit Models of Language, Vision, and Action, Saarbrücken, Germany {yifwang,mayank,vera}@lst.uni-saarland.de, sukrut.rao@mpi-inf.mpg.de, ji-ung.lee@uni-saarland.de
@@ 1 Introduction
Pre-trained language models (PLMs) such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) have significantly advanced performance across a plethora of NLP tasks (Wang et al., 2018; Gao et al., 2023). However, their complex architectures and black-box nature make understanding their behavior a persistent challenge (Bommasani et al., 2021). To address this, research has increasingly focused on explaining model predictions, particularly in relation to the input. These input-based explanations, often referred to as local explanations or rationales, aim to reveal how specific inputs influence a model’s predictions (Arras et al., 2019; Atanasova et al., 2020; Lyu et al., 2024).
@@
Figure 1: Visualization of $\mathbf{W(x)x}$ in a conventionally fine-tuned model (Conventional LM) and a B-cos LM. Green (red) indicates the positive (negative) impact of tokens on the prediction. In both examples, both models correctly predict not toxic. In the Conventional LM, “funny” is incorrectly assigned a negative attribution in example (a), while in example (b), irrelevant words like “why” and “smell” are highlighted, making the explanations unfaithful and less interpretable.
@@
Most explanation methods for neural models are post-hoc, meaning that they attempt to explain a model’s behavior only after it has been trained and deployed (Sundararajan et al., 2017; Ribeiro et al., 2016). While these methods are widely used and easy to apply, they have been shown to produce unfaithful and less interpretable explanations (Smilkov et al., 2017; Kindermans et al., 2019; Slack et al., 2020; Pruthi et al., 2020). [^1] Considering the evolving definition of these terms in past literature, we provide a detailed definition in Appendix A. Prior research has attributed these shortcomings to the lack of explainability in contemporary neural models (Kindermans et al., 2018; Alvarez Melis and Jaakkola, 2018; Rudin, 2019). Figure 1 provides examples illustrating this issue.
@@
To overcome these limitations, we introduce B-cos LM, a dynamic linear model that learns task-relevant patterns through increased input-weight alignment pressure. Building upon B-cos networks from computer vision (Böhle et al., 2022; Arya et al., 2024), we improve explainability of B-cos LMs through mathematically grounded architectural and computational adaptations. Furthermore, we tailor B-cos LMs for NLP by incorporating specialized architectural modifications and training pipelines. Our contributions are as follows:
@@
    Based on our findings, we provide practical guidelines for building effective B-cos LMs.
@@ 2 Related Work
Various methods have been proposed to provide post-hoc explanations for neural model predictions (Atanasova et al., 2020). These methods can be broadly categorized based on how they generate explanations: gradient-based (Simonyan et al., 2014; Kindermans et al., 2016; Sundararajan et al., 2017), propagation-based (Bach et al., 2015; Shrikumar et al., 2017; Springenberg et al., 2015), and perturbation-based methods (Li et al., 2016; Ribeiro et al., 2016; Lundberg and Lee, 2017). Besides, the attention mechanism (Bahdanau et al., 2015) is often viewed as an explanation, particularly in transformer-based models (Vaswani et al., 2017).
Although post-hoc methods can be applied to generate explanations for existing models, numerous studies have shown that they lack faithfulness, often failing to capture the true decision-making process of the model (Kindermans et al., 2019; Jain and Wallace, 2019; Slack et al., 2020; Pruthi et al., 2020). Furthermore, they may generate noisy explanations that focus on irrelevant information, making them difficult for humans to interpret (Smilkov et al., 2017; Ismail et al., 2021).
@@
The limitations of post-hoc explanation methods may be attributed to the inherent lack of explainability in contemporary neural models, which are typically optimized solely for task performance (Kindermans et al., 2018; Rudin, 2019; Atanasova et al., 2022). For instance, studies have shown that existing models struggle to provide faithful explanations (Alvarez Melis and Jaakkola, 2018) or tend to learn noisy patterns, resulting in less interpretable explanations (Ismail et al., 2021).
In response, various efforts have been made to enhance model explainability. Some work has introduced constraints that improve specific explanation properties, such as faithfulness (Tutek and Šnajder, 2022; Moradi et al., 2020, 2021), consistency (Atanasova et al., 2022), locality (Alvarez Melis and Jaakkola, 2018), and plausibility (Ismail et al., 2021). However, as these constraints are typically imposed as regularizers, their effectiveness in improving explanation quality is not guaranteed (Pruthi et al., 2020). Others have proposed self-explanatory model architectures such as rationale-based models that utilize an “explain-then-predict” pipeline, where one module selects rationales for another to make predictions based on them (Lei et al., 2016). Although seemingly transparent, both components rely on neural networks, making the rationale extraction and utilization processes opaque (Zheng et al., 2022; Jacovi and Goldberg, 2021). Besides, such models may face optimization challenges that limit their practicality in real-world tasks (Lyu et al., 2024).
@@
Property Conventional Fine-tuning B-cosification (Arya et al., 2024) B-cos LM (ours)
@@
Table 1: Comparison between conventional fine-tuning, B-cosification in computer vision and B-cosification in NLP (B-cos LM). Conventional fine-tuning and B-cosification follow the configuration of BERT for sequence classification and CLIP (Radford et al., 2021), respectively (cf. § 3 for details).
3 Methodology
In this section, we outline the architecture and training process of B-cos LMs and how their design ensures faithful and human interpretable explanations. We first introduce B-cos networks (§ 3.1) and then describe how we transform PLMs to task-specific B-cos LMs (§ 3.2). Finally, we demonstrate how to generate explanations from B-cos LMs (§ 3.3). Notations used in the work are detailed in Appendix B.
3.1 B-cos Networks
Complex neural networks can be interpreted as generalized linear models (Nair and Hinton, 2010; Alvarez Melis and Jaakkola, 2018; Srinivas and Fleuret, 2019). For each input $\mathbf{x}$, the network applies a linear transformation: $\mathbf{f(x)=W(x)x+b(x)}$, where both the weight $\mathbf{W(x)}$ and bias $\mathbf{b(x)}$ depend on $\mathbf{x}$. Given that many activation functions are (approximately) piecewise linear, the overall network can be viewed as (approximately) piecewise affine (Alvarez Melis and Jaakkola, 2018). Earlier work refers to such models as dynamic linear models (Böhle et al., 2021, 2022), highlighting the fact that the weight and bias terms dynamically change according to $\mathbf{x}$.
Under this dynamic linear perspective, the linear mapping $\mathbf{W(x)}$ can be seen as attributing model predictions to individual input features. However, two challenges hinder the direct use of this interpretation. First, $\mathbf{W(x)}$ alone provides an incomplete and unfaithful explanation since $\mathbf{f(x)\neq W(x)x}$ due to the presence of the bias term $\mathbf{b(x)}$, and incorporating $\mathbf{b(x)}$ into explanations is highly non-trivial (Wang et al., 2019). Second, $\mathbf{W(x)}$ is often difficult for humans to interpret, as it does not necessarily align only with task-relevant input patterns (Smilkov et al., 2017) and therefore yields noisy and irrelevant explanations. Figure 1 illustrates these challenges. To address these issues, Böhle et al. (2022) introduced B-cos networks by replacing the conventional linear transformation:
with a B-cos transformation:
B-cos $\mathbf{(x;w)}=\mathbf{\hat{w}^{T}x}\times|\text{cos}\mathbf{(x,\hat{w})|^{B-1}}$ (2)
$=\|\mathbf{\hat{w}\|\|x\|}|\text{cos}\mathbf{(x,\hat{w})|^{B}}\times\mathbf{\text{sgn}(\text{cos}(x,\hat{w}))}$
B-cos$\mathbf{(x;w)}$ can be seen as a linear transformation of $\mathbf{x}$ with the dynamic linear weight $\mathbf{w(x)=|\text{cos}(x,\hat{w})|^{B-1}\times\hat{w}}$. The absence of $\mathbf{b(x)}$ ensures the completeness of summary $\mathbf{w(x)}$. We demonstrate that this completeness extends to an entire network composed of bias-free dynamic linear modules in 3.3. Moreover, with additional alignment pressure (B>1), the weight $\mathbf{w}$ is forced to align closely with task-relevant patterns to achieve a high cosine similarity and strong activation within the B-cos module. As a result, only the most relevant features are highlighted in explanations, making them more interpretable to humans.
While early B-cos models were trained from scratch, Arya et al. (2024) recently introduced B-cosification, an efficient method to obtain B-cos models. This approach first modifies conventional models with task capacities to adopt the B-cos architecture, followed by fine-tuning on downstream datasets for B-cos conversion. B-cosified models generate explanations as faithful and interpretable as B-cos models trained from scratch but at a much lower training cost. However, directly applying B-cosification to NLP models is non-trivial and inefficient due to the significant differences in model architectures and training pipelines.
3.2 B-cosification in NLP
In this section, we present our B-cosification approach for NLP. We summarize the differences between B-cosification in NLP, its counterpart in vision, and conventional fine-tuning in Table 1. We provide an extensive ablation study in Appendix C.
@@ 3.2.1 B-cos Adaptations
For completeness and faithfulness of explanations, we follow Arya et al. (2024) and remove all bias terms in models, including those in the affine transformations of layer normalization and attention blocks. Additionally, a prediction head is typically added on top of the transformer before fine-tuning for downstream tasks in the NLP pipeline. This head often includes activation functions that are not (approximately) piecewise linear, such as sigmoid and tanh. To accommodate the unique architecture of NLP models, we remove all activation functions in the prediction heads, as they reduce the locality of explanations and introduce numerical instability during their generation. We expect the added non-linearity from B>1 to compensates for this removal.
@@
To promote input-weight alignment and improve human interpretability of explanations, we follow Arya et al. (2024) and replace all linear transformations with B-cos transformations in § 3.1. For a more efficient B-cosification, B-cos layers are initialized with the corresponding weights $\mathbf{W}$ of the original model.

### v2

@@ Abstract
Post-hoc explanation methods for black-box models often struggle with faithfulness and human interpretability due to the lack of explainability in current neural architectures. Meanwhile, B-cos networks have been introduced to improve model explainability by proposing an architecture that removes bias terms and promotes input-weight alignment. Although B-cos networks have shown success in building explainable systems, their application has so far been limited to computer vision models and their associated training pipelines. In this work, we introduce B-cos LMs, i.e., B-cos language models (LMs) empowered for natural language processing (NLP) tasks. Our approach directly transforms pre-trained language models into B-cos LMs by combining B-cos conversion and task fine-tuning, improving efficiency compared to previous methods. Our automatic and human evaluation results demonstrate that B-cos LMs produce more faithful and human interpretable explanations than post-hoc methods, while maintaining task performance comparable to conventional fine-tuning. Our in-depth analysis explores how B-cos LMs differ from conventionally fine-tuned models in their learning processes and explanation patterns. Finally, we are also the first to explore the transformation of decoder-only models to B-cos LMs for generation tasks.
@@ 1 Introduction
Pre-trained language models (PLMs) such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) have significantly advanced performance across a plethora of NLP tasks (Wang et al., 2018; Gao et al., 2023). However, their complex architectures and black-box nature make understanding their behavior a persistent challenge (Bommasani et al., 2021). To address this, research has increasingly focused on understanding model predictions in various natural language understanding and generation tasks using different forms of explanations, such as input-based explanations (Feng et al., 2024; Wei Jie et al., 2024; Jiang et al., 2024; Madsen et al., 2024; Yin & Neubig, 2022; Deiseroth et al., 2023), natural language explanations (Ramnath et al., 2024; Wang et al., 2025), and concept-based explanations (Yu et al., 2024; Raman et al., 2024). Among others, input-based explanations, often referred to as rationales, aim to reveal how specific inputs influence a model’s prediction (Arras et al., 2019; Atanasova et al., 2020; Lyu et al., 2024). In this work, we focus on input-based explanations, as they offer the most direct insight into model behavior and are often mandated by laws, such as the EU Artificial Intelligence Act.
@@
Figure 1: Visualization of $\mathbf{W(x)x}$ in a conventionally fine-tuned model (Conventional LM) and a B-cos LM. Green (red) indicates the positive (negative) impact of tokens on the prediction. In both examples, both models correctly predict not toxic. In the Conventional LM, “funny” is incorrectly assigned a negative attribution in example (a), and in example (b), irrelevant words like “why” and “smell” are highlighted, making the explanations unfaithful and less interpretable. Examples and explanations are drawn from HateXplain. See  §3 for details on how $\mathbf{W(x)x}$ is computed.
@@
Most input-based explanation methods for neural models are post-hoc, meaning that they attempt to explain a model’s behavior only after it has been trained and deployed (Sundararajan et al., 2017; Ribeiro et al., 2016). While these methods are widely used and easy to apply, they have been shown to produce unfaithful explanations that do not accurately reflect the model’s actual reasoning process (Kindermans et al., 2019; Slack et al., 2020; Pruthi et al., 2020). They also struggle with human interpretability, making it difficult for users to understand the model’s reasoning (Smilkov et al., 2017; Ismail et al., 2021). Prior research suggests that these limitations stem from a lack of inherent explainability in current models, that is, the model’s ability to generate faithful and interpretable explanations by design (Kindermans et al., 2018; Alvarez Melis & Jaakkola, 2018; Rudin, 2019). As a result, improving model explainability is crucial for producing explanations that are both reliable and useful to users. [^1] Considering the evolving definition of these terms in past literature, we provide a detailed definition in Appendix A. Figure 1 provides examples illustrating this issue.
@@
To overcome these limitations, we introduce B-cos LM, a dynamic linear model that learns the most task-relevant patterns through increased input-weight alignment pressure. Building upon B-cos networks that were first introduced by Böhle et al. (2022) for computer vision, we ensure the explainability of B-cos LMs through mathematically grounded architectural and computational adaptations, with specialized architectural modifications and training pipelines tailored for NLP tasks.
We conduct comprehensive empirical experiments using encoder-only models on classification tasks. Our focus on classification is motivated by its prevalence in high-stakes applications, such as loan approvals, hiring decisions, and hate speech detection, where explainability is crucial and often legally mandated. Encoder-only models have also seen renewed interest in the research community (Warner et al., 2024; Breton et al., 2025; Chaffin, 2025), and they remain the standard architecture for text classification and continue to perform competitively compared to large language models (LLMs) (Zhao et al., 2024). Beyond that, we also explore applying B-cos LMs to decoder-only models for generation tasks and show that B-cos LMs can be extended to a variety of tasks and the latest model architectures. Our contributions are as follows:
@@
    We present a first study of applying B-cos LMs to decoder-only models and generation tasks, providing a step towards a broader application of B-cos LMs in the era of LLMs.
@@ 2 Related Work
Various methods have been proposed to provide post-hoc explanations for neural model predictions (Atanasova et al., 2020). These methods can be broadly categorized based on how they generate explanations: gradient-based (Simonyan et al., 2014; Kindermans et al., 2016; Sundararajan et al., 2017; Enguehard, 2023), propagation-based (Bach et al., 2015; Shrikumar et al., 2017; Springenberg et al., 2015; Ferrando et al., 2023; Modarressi et al., 2022; 2023), and perturbation-based methods (Li et al., 2016; Ribeiro et al., 2016; Lundberg & Lee, 2017; Deiseroth et al., 2023). Besides, the attention mechanism (Bahdanau et al., 2015) is often viewed as an explanation, particularly in transformer-based models (Vaswani et al., 2017). While most existing work focuses on understanding model predictions in classification settings, recent efforts have also aimed to explain model behavior in generation tasks, including sentence completion (Yin & Neubig, 2022; Ferrando et al., 2023), question answering (Enouen et al., 2024), and summarization (Cohen-Wang et al., 2024).
Although post-hoc methods have been widely used, numerous studies have shown that they lack faithfulness, often failing to capture the true decision-making process of the model (Kindermans et al., 2019; Jain & Wallace, 2019; Slack et al., 2020; Pruthi et al., 2020). Furthermore, they are noisy and may select irrelevant information leading to explanations that cannot be interpreted by humans (Smilkov et al., 2017; Ismail et al., 2021).
@@
Prior research suggests that the lack of faithfulness and human interpretability in post-hoc explanations arises from the fundamental lack of explainability in modern neural models, which are typically optimized solely for task performance (Kindermans et al., 2018; Rudin, 2019; Atanasova et al., 2022). In response, various efforts have been made to enhance model explainability. Some works have introduced constraints that improve specific explanation properties, such as faithfulness (Tutek & Šnajder, 2022; Moradi et al., 2020; 2021; Barkan et al., 2024), consistency (Atanasova et al., 2022), locality (Alvarez Melis & Jaakkola, 2018), and plausibility (Ismail et al., 2021). However, as these constraints are typically imposed as regularizers, their effectiveness in improving explanation quality is not guaranteed (Pruthi et al., 2020). Others have proposed self-explanatory model architectures such as rationale-based models that utilize an “explain-then-predict” pipeline, where one module selects rationales for another to make predictions based on them (Lei et al., 2016). Although seemingly transparent, both components rely on neural networks, making the rationale extraction and utilization processes opaque (Zheng et al., 2022; Jacovi & Goldberg, 2021). Besides, such models may face optimization challenges that limit their practicality in real-world tasks (Lyu et al., 2024).
@@
3 Methodology
In this section, we outline the architecture and training process of B-cos LMs and how their design ensures faithful and human interpretable explanations. We first introduce B-cos networks (§ 3.1 and § 3.2) and then describe how we transform PLMs to task-specific B-cos LMs (§ 3.3). Finally, we demonstrate how to generate explanations from B-cos LMs (§ 3.4). Notations used in the work are detailed in Appendix B.
3.1 B-cos Networks
Complex neural networks can be interpreted as generalized linear models (Nair & Hinton, 2010; Alvarez Melis & Jaakkola, 2018; Srinivas & Fleuret, 2019). For each input $\mathbf{x}$, the network effectively applies a linear transformation: $\mathbf{f(x)=W(x)x+b(x)}$, where both the weight $\mathbf{W(x)}$ and bias $\mathbf{b(x)}$ depend on $\mathbf{x}$. Given that many activation functions are (approximately) piecewise linear, the overall network can be viewed as (approximately) piecewise affine (Alvarez Melis & Jaakkola, 2018). Earlier work refers to such models as dynamic linear models (Böhle et al., 2021; 2022), highlighting the fact that the weight and bias terms dynamically change according to $\mathbf{x}$.
Under this dynamic linear perspective, the linear mapping $\mathbf{W(x)}$ can be seen as attributing model predictions to individual input features, and $\mathbf{W(x_{i})x_{i}}$ can be seen as the contribution of feature $\mathbf{x_{i}}$ to the model prediction. However, two challenges hinder the direct use of this interpretation. First, $\mathbf{W(x)}$ alone provides an incomplete and unfaithful model summary since $\mathbf{f(x)\neq W(x)x}$ due to the presence of the bias term $\mathbf{b(x)}$, and incorporating $\mathbf{b(x)}$ into explanations is highly non-trivial (Wang et al., 2019). Second, $\mathbf{W(x_{i})x_{i}}$ is often difficult for humans to interpret, as $\mathbf{W(x)}$ does not necessarily align only with task-relevant input patterns (Smilkov et al., 2017) and therefore yields noisy and irrelevant explanations. Figure 1 illustrates these challenges. To address these issues, Böhle et al. (2022) introduced B-cos networks by replacing the conventional linear transformation:
with a B-cos transformation:
B-cos $\mathbf{(x;w)}=\mathbf{\hat{w}^{T}x}\times|\text{cos}\mathbf{(x,\hat{w})|^{\text{B-1}}}$ (2)
$=\|\mathbf{\hat{w}\|\|x\|}|\text{cos}\mathbf{(x,\hat{w})|^{\text{B}}}\times\mathbf{\text{sgn}(\text{cos}(x,\hat{w}))}$
B-cos$\mathbf{(x;w)}$ can be seen as a linear transformation of $\mathbf{x}$ with the dynamic linear weight $\mathbf{w(x)=|\text{cos}(x,\hat{w})|^{\text{B-1}}\times\hat{w}}$. The absence of $\mathbf{b(x)}$ ensures the completeness of summary $\mathbf{w(x)}$. We demonstrate that this completeness extends to an entire network composed of bias-free, dynamic linear modules in § 3.4. Moreover, since the B-cos module output is bounded by $\mathbf{\|x\|}$, the weight $\mathbf{w}$ must align closely with task-relevant patterns to achieve a high cosine similarity and strong activation, especially under additional alignment pressure (B>1). This drives the model to assign greater weight to the most relevant features when optimizing target output probabilities, promoting the learning of representative patterns during training. Consequently, during explanation generation, task-relevant features $\mathbf{x_{i}}$ receive higher attribution $\mathbf{W(x_{i})x_{i}}$ due to stronger alignment, while irrelevant features receive lower attribution, suppressed by weaker alignment and the exponential scaling. For a more detailed discussion of how the B-cos transformation enhances faithfulness and human interpretability, see Böhle et al. (2022; 2024).
While early B-cos models were trained from scratch, Arya et al. (2024) recently introduced B-cosification, an efficient method to obtain B-cos models. This approach first modifies conventional models with task capacities to adopt the B-cos architecture, followed by fine-tuning on downstream datasets for B-cos conversion. B-cosified models generate explanations as faithful and interpretable as B-cos models trained from scratch but at a much lower training cost. However, directly applying B-cosification to LMs is non-trivial and inefficient due to the significant differences in model architectures and training pipelines.
3.2 Dynamic Linear Representation of Model Components
Here we describe how each model component in transformers can function as or be converted to a bias-free, dynamic linear module in B-cos LMs.
B-cos Layers
[…]
Non-linear Activation Functions
In transformer models, non-linearity is typically introduced using (approximately) piecewise linear activation functions, such as ReLU (Nair & Hinton, 2010) and GELU (Hendrycks & Gimpel, 2016). These functions can be easily interpreted as linear transformations with input-dependent weights. For example, $\text{GELU}(\mathbf{x})=\mathbf{x}\times(0.5+0.5\times\text{erf}(\mathbf{x}/\sqrt{2}))$ can be interpreted as a linear transformation where the second term acts as the dynamic linear weight.
Attention Blocks
$\text{Att}(\mathbf{X;Q,}\mathbf{K,V})=\text{softmax}(\mathbf{X^{T}Q^{T}KX)VX}=\mathbf{A(X)VX}=\mathbf{W(X)X}$ (3)
$\text{MSA}\mathbf{(X)=U[W}_{1}\mathbf{(X)X,...,W}_{H}\mathbf{(X)X)]}$ (4)
Normalization Layers
Following Böhle et al. (2024), we apply bias-free normalization layers to ensure completeness of explanations:
$\star\mathbf{Norm(x,\mathcal{X};\gamma)}=\frac{\mathbf{x-\langle\mathcal{X}\rangle_{\star}}}{\sqrt{\text{var}_{\star}\mathbf{(\mathcal{X})}}}\times\mathbf{\gamma}$ (5)
[…]
3.3 B-cosification for LMs
In this section, we present our B-cosification approach for LMs. We summarize the differences between B-cosification for LMs, its counterpart for vision models, and conventional fine-tuning in Table 1.
Property Conventional Fine-tuning B-cosification for vision (Arya et al., 2024) B-cos LM (ours)
@@ 3.3 B-cosification for LMs
Table 1: Comparison between conventional fine-tuning, B-cosification for computer vision models and B-cosification for language models (B-cos LM). Conventional fine-tuning and B-cosification for vision follow the configuration of BERT for sequence classification and CLIP (Radford et al., 2021), respectively (cf. § 3 for details).
@@ 3.3.1 B-cos Adaptations
For completeness and faithfulness of explanations, we follow Arya et al. (2024) and remove all bias terms in models, including those in the affine transformations of layer normalization and attention blocks. Additionally, a prediction head is typically added on top of the transformer before fine-tuning for downstream tasks in the NLP pipeline. This head often includes activation functions that are not (approximately) piecewise linear, such as sigmoid and tanh. To accommodate the unique architecture of LMs, we remove all activation functions in the prediction heads, as they generate explanations that are not locally difference-bounded (Alvarez Melis & Jaakkola, 2018) and introduce numerical instability during explanation generation. Our experiments show that the added non-linearity from B>1 could compensate for this removal.
@@
To promote input-weight alignment and improve human interpretability of explanations, we replace all linear transformations with B-cos transformations in § 3.1. For a more efficient B-cosification, B-cos layers are initialized with the corresponding weights of the original model.